OpenStudy (anonymous):

help me PWEASEEEEE In order to understand the prehistory of the Hawaiian island of Lana'i better, anthropologists Maria Sweeney, Melinda Allen, and Boyd Dixon used radiocarbon dating on charcoal found in an ancient dwelling site, the Kaunolu Village National Historic Landmark, the largest archeological complex on the island. In one of their samples, they found that approximately 94% of the original carbon 14 remained. Using the fact that Carbon 14 decays by 1.202% every 100 years, determine the approximate age of the this sample .

OpenStudy (anonymous):

1200 years

OpenStudy (anonymous):

... how do u know i dont need the answer i need to know how...

OpenStudy (anonymous):

The voice in my head told me.

OpenStudy (anonymous):

you need a formula for this

OpenStudy (anonymous):

im guessing so ... i think it involves LOGs

OpenStudy (anonymous):

decays be 1.202% every 100 years means every 100 years you would multiply by 100%-1.202%= .98798

OpenStudy (anonymous):

so you can use \[A=A_0\times (.98798)^{\frac{t}{100}}\]

OpenStudy (anonymous):

what is the 0 after A signify?

OpenStudy (anonymous):

don't worry about the \(A\) and the \(A_0\)

OpenStudy (anonymous):

A is your amount in \(t\) years, it is a function \(A(t)\) and \(A_0\) is what you start with

OpenStudy (anonymous):

oh ok so logarithms aren't used in this problem?

OpenStudy (anonymous):

but you don't need to worry about that, because whatever you start with you know you have 94% left, so solve \[.94=(.98798)^{\frac{t}{100}}\] for \(t\) and yes, you need logs

OpenStudy (anonymous):

oh ok give me a sec i wanna try this on paper

OpenStudy (anonymous):

ok should take two steps only

OpenStudy (anonymous):

ok im lost i got log 94/100/ log(1-.01202)

OpenStudy (anonymous):

did you start with \[ .94=(.98798)^{\frac{t}{100}}\]?

OpenStudy (anonymous):

oh looks sort of good but you have too many fraction bars

OpenStudy (anonymous):

the log94/100 is over log(1-.1202)

OpenStudy (anonymous):

\[\frac{t}{100}=\frac{\ln(.94)}{\ln(.98798)}\] \[t=100\times \frac{\ln(.94)}{\ln(.98798)}\]

OpenStudy (anonymous):

wait u lost me im supposed to do this on excel and i have no clue how that applies

OpenStudy (anonymous):

me neither i used this \[A=b^x\iff x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

hmmm ok im gonna try to figure this out by myself give me a couple

OpenStudy (anonymous):

was the answer 5.12

OpenStudy (anonymous):

off by some decimal places http://www.wolframalpha.com/input/?i=100ln%28.94%29%2Fln%28.98798%29

OpenStudy (anonymous):

it is 5.12 before you multiply by 100

OpenStudy (anonymous):

because the exponent is not \(t\) it is \(\frac{t}{100}\) so you have to multiply by 100 to get \(t\) after taking the log